Circular No. 6049

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V1974 CYGNI L. Rosino, P. Rafanelli, T. Iijima, and E. Esenoglu, Asiago Astrophys-

ical Observatory, report: “CCD spectra of the nova V1974 Cyg, obtained in June and July with the 1.82-m Asiago telescope at Mt. Ekar, indicate a further decrease of the degree of ionization. The strongest emission lines observed in the optical spectrum and their approximate intensities relative to Hβ (absolute flux 31.5 × 10^{−14} erg cm^{−2} s^{−1}) are now, in order of decreasing strength: [Ne V] at 342.6 nm, 19.9; [O III] at 500.7 nm, 17.8; Hα (possibly blended with [N II]), 5.2; [Ne III] at 386.9 nm, 2.8; He II at 468.6 nm, 1.2. Other emission lines, with intensities from 0.55 to 0.1, observed in the spectra are: Hδ + N III; Hγ; [O III] at 436.3 nm; He I at 587.6, 667.8, and 706.5 nm; N III at 464.0 nm; [Fe VI] at 567.7 nm; [Fe VII] at 608.7 nm; [O I] at 630.0 nm; and [O II] at 732–733 nm. No further traces have been found of the coronal lines observed in 1993. The approximate FWHM expansion velocity of the ejecta is of the order of 1600 ± 100 km/s. The strongest emission lines in the spectrum display typical saddle-shaped profiles.”

V407 CYGNI U. Munari, Asiago Astrophysical Observatory; A. Bragaglia and M.

D. Guarnieri, Bologna Astronomical Observatory (BAO); G. Sostero and A. Lepardo, Associazione Friulana Astronomia e Meteorologia; and B. F. Yudin, Moscow Astronomical Institute (MAI), communicate: “The symbiotic Mira-type variable V407 Cyg (= Nova Cyg 1936) is currently undergoing an unexpected bright and active phase (cf. Munari et al. 1990, MNRAS 242, 653). Photoelectric photometry secured with the 0.91-m telescope of the Catania Astronomical Observatory on Aug. 4 and 6 gives V = 13.0, B − V = +1.0, U − B = 0.0. CCD spectra (range 390-780 nm; 1-nm resolution) obtained with the 1.5-m BAO telescope (+ BFOSC) on Aug. 7.0 UT show the Mira absorption continuum overwhelmed by a hot continuum shortward of 540 nm. Hydrogen, He I, and forbidden O III lines (never seen in the Asiago 1986–1993 low-resolution spectra) are in strong emission. JHKL photometry secured in July with the 1.25-m MAI Crimean telescope confirms that the Mira variable is now close to minimum brightness during its 745-day pulsation cycle.”

1994 August 8 Daniel W. E. Green